the time in seconds required to produce a P.D at 20v across a capacitor at 1000uF when it is charged at a steady rate of 200uC/sec?

Grade:12

2 Answers

Abhaya kumar mishra abhaya
11 years ago

100 secs.

Srikanth Neeli
11 years ago

potential difference=20v,capacity=1000uF=1/1000F,I=200uC/s=0.2/1000C/s

 

from capacitance Q=CV,

                 we get,Q=0.02C

 

from current definition, t=Q/i

                                  t=0.02*1000/0.2

                                  t=100sec

 

 

IF YOU SATISFY the answer PLEASE CLICK ON "YES",please dont forget...............